# Existence, Sharp Boundary Asymptotics, and Stochastic Optimal Control for Semilinear Elliptic Equations with Gradient-Dependent Terms and Singular Weights

### 08 Mar 2026 v1

@@ 2.1. The Defining Function
The strict convexity of $\Omega$ (Assumption 1.1) implies the existence of a smooth strictly convex defining function. Following the approach of Gilbarg and Trudinger [13] and Lazer and McKenna [19], we have:
@@
There exists $\varphi\in C^{2}(\overline{\Omega})$ satisfying:
@@
    $D^{2}\varphi(x)\geq\kappa_{0}I$ for all $x\in\overline{\Omega}$, where $\kappa_{0}>0$ is the minimum principal curvature of $\partial\Omega$.
@@
For $x$ sufficiently close to $\partial\Omega$, set $\varphi(x)=-d(x)$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. The signed distance function is $C^{2}$ in a tubular neighborhood of $\partial\Omega$ when the boundary is $C^{2,\alpha}$. The Hessian of d-d equals the second fundamental form of the level set $\{d=\text{const}\}$, which is positive definite by strict convexity. The function can be extended to all of $\overline{\Omega}$ while preserving strict convexity; see [13, Chapter 14]. ∎
Let $v(x)=-\varphi(x)$. Then $v>0$ in $\Omega$, $v=0$ on $\partial\Omega$, and $v$ is strictly concave ($D^{2}v<0$). By construction, $v(x)=d(x)$ identically in a tubular neighborhood of $\partial\Omega$. Define
$m_{0}=\min_{x\in\partial\Omega}|\nabla v(x)|=1,\quad m_{1}=\max_{x\in\partial\Omega}|\nabla v(x)|=1.$ (2.1)
By continuity, for any $\epsilon>0$, there exists $\delta>0$ such that $1-\epsilon\leq|\nabla v(x)|\leq 1+\epsilon$ whenever $d(x)<\delta$.
@@ 2.2. Explicit Barrier Construction
We now construct explicit sub- and supersolutions that determine the precise boundary blow-up rate.
@@
    Critical Logarithmic Case ($q=\beta+2$): The singular growth is logarithmic, $u(x)\sim C\ln(1/v(x))$. There exist $C_{+}>C_{-}^{*}>0$ and $0<C_{-}<C_{-}^{*}$ providing local barriers near boundary.
@@
We provide detailed computations for each case.
@@
We compute:
@@
Step 2: Substitution into the operator.
Substituting (2.4)–(2.6) into $\mathcal{L}W$:
@@
$\quad+b(x)h(C\gamma v^{-\gamma-1}|\nabla v|)+a(x)Cv^{-\gamma}-f(x).$ (2.7)
@@
Using the rigorous exact bounds derived from (1.2)–(1.3), for any $\varepsilon>0$, there exists $\delta>0$ such that for all $x\in\Omega$ with $v(x)<\delta$, we have
$(b_{1}-\varepsilon)v(x)^{\beta}\leq b(x)\leq(b_{2}+\varepsilon)v(x)^{\beta}.$
Additionally, since $v(x)=d(x)$ near $\partial\Omega$, $|\nabla v(x)|=1$. The analysis for the supersolution exact requirement proceeds as:
@@
$=(b_{1}-\varepsilon)l_{1}(C\gamma)^{q}v^{\beta-q(\gamma+1)}.$
The mathematically exact singular terms determining the sign of $\mathcal{L}W$ are isolated as:
@@
where $\mathcal{R}(x)$ denotes terms of strictly lower singular order relative to $v^{-\gamma-2}$.
We choose $\gamma$ to perform an exact balance of the leading singularities by setting:
$\beta-q(\gamma+1)+\gamma+2=0\iff\gamma(1-q)=\beta-q+2\iff\gamma=\frac{\beta-q+2}{q-1}.$
Note that $\gamma>0$ iff $\beta-q+2>0$ iff $q<\beta+2$.
With this choice, ensuring supersolution positivity reduces to enforcing the strict inequality:
$\liminf_{x\rightarrow\partial\Omega}\left(v^{\gamma+2}\mathcal{L}W\right)\geq(b_{1}-\varepsilon)l_{1}\gamma^{q}C^{q}-\gamma(\gamma+1)C>0.$ (2.9)
Define
$P(C)=b_{1}l_{1}\gamma^{q}C^{q}-\gamma(\gamma+1)C.$
@@
  • $P(C)<0$ for small $C>0$ (linear term dominates);
  • $P(C)>0$ for large $C$ (power term dominates);
  • $P(C^{*})=0$ where $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}l_{1}\gamma^{q}}\right)^{1/(q-1)}$.
@@
Step 4: Verification of the subordinate terms.
The reaction term rigorous limit provides the bound
$a(x)Cv^{-\gamma}\leq a_{2}Cv^{\alpha-\gamma}.$
Since we require $\alpha>-2$ (Assumption 1.4), we explicitly have
$\alpha-\gamma>-2-\gamma=-\gamma-2,$
so the reaction term is of lower singular order than the leading terms and does not affect the balance.
@@
Step 5: Case 2 – High-order gradient.
When $q>\beta+2$, we have $\beta-q(\gamma+1)<-\gamma-2$ for any $\gamma>0$. Thus the gradient term $v^{\beta-q(\gamma+1)}$ is more singular than the Laplacian term $v^{-\gamma-2}$. The gradient term always dominates, making $\mathcal{L}W>0$ near $\partial\Omega$ for any $C>0$.
Step 6: Case 3 – Critical logarithmic.
When $q=\beta+2$, let
$W=C\ln(1/v)=-C\ln v.$
Then:
$\nabla W$ $=-Cv^{-1}\nabla v,$
$|\nabla W|$ $=Cv^{-1}|\nabla v|,$
$\Delta W$ $=Cv^{-2}|\nabla v|^{2}-Cv^{-1}\Delta v.$
Substituting:
$\mathcal{L}W$ $\geq-Cv^{-2}|\nabla v|^{2}+Cv^{-1}\Delta v+(b_{1}-\varepsilon)v^{\beta}l_{1}(Cv^{-1})^{q}|\nabla v|^{q}+a_{1}v^{\alpha}C\ln(1/v)-f$
$=v^{-2}\left[(b_{1}-\varepsilon)l_{1}C^{q}v^{\beta-q+2}-C\right]+\mathcal{R}_{L}(x),$
where $\mathcal{R}_{L}(x)$ denotes terms of strictly lower singular order relative to $v^{-2}$ near $\partial\Omega$. Since $\beta-q+2=0$ in the exact critical case, we rigorously obtain
$\liminf_{x\rightarrow\partial\Omega}\left(v^{2}\mathcal{L}W\right)\geq(b_{1}-\varepsilon)l_{1}C^{q}-C.$
The same argument as in Step 3 shows that barriers exist with the appropriate choice of $C_{\pm}$. ∎
@@
The strict convexity of $\Omega$ is essential in two ways: (i) it guarantees the existence of a smooth, strictly concave defining function $v$, ensuring $D^{2}v<0$; (ii) it ensures that $|\nabla v|$ is bounded away from zero near $\partial\Omega$, which is necessary for the gradient term to provide the required singular behavior.
@@ 3. Existence, Uniqueness, and Convexity
$u$ $\leq w\quad\text{on }\partial\Omega^{\prime}.$
@@
for some $x_{0}\in\Omega^{\prime}$. At this interior maximum point:
$\nabla u(x_{0})$ $=\nabla w(x_{0}),$
$\Delta u(x_{0})$ $\leq\Delta w(x_{0}).$
Therefore,
@@
$=-\Delta u(x_{0})+\Delta w(x_{0})+b(x_{0})[h(|\nabla u(x_{0})|)-h(|\nabla w(x_{0})|)]$
$\quad+a(x_{0})[u(x_{0})-w(x_{0})]$
$\geq a(x_{0})[u(x_{0})-w(x_{0})]>0,$
a contradiction. Here we used that $h(|\nabla u|)=h(|\nabla w|)$ at $x_{0}$ since the gradients coincide. ∎
@@
Let $\overline{u}(x)=C_{+}v(x)^{-\gamma}$ and $\underline{u}(x)=C_{-}v(x)^{-\gamma}$ be the super- and subsolutions from Lemma $2.$ By construction, these satisfy
@@
for some $\delta_{0}>0$.
For the interior region
$\Omega_{\delta_{0}}=\{x\in\Omega:d(x)\geq\delta_{0}\},$
we solve the Dirichlet problem
@@
Similarly construct a global subsolution $\underline{U}$ with $\underline{U}\leq\overline{U}$.
@@
$\mathcal{S}=\left\{w\in C(\Omega):w\text{ is a subsolution},\underline{U}\leq w\leq\overline{U}\right\}.$
@@
Taking the supremum over all such modifications, standard arguments [10] show that $u$ is a viscosity solution. By interior regularity for uniformly elliptic equations with Hölder continuous coefficients, $u\in C^{2,\alpha}_{\mathrm{loc}}(\Omega)$, and $u$ is a classical solution.
@@
Suppose $u_{1}$ and $u_{2}$ are two solutions satisfying (1.4). For any $\epsilon>0$, define
$\Omega_{\epsilon}=\{x\in\Omega:d(x)>\epsilon\}.$
On $\partial\Omega_{\epsilon}$:
$|u_{1}(x)-u_{2}(x)|\leq(C_{+}-C_{-})\epsilon^{-\gamma}.$
Since both $u_{1}$ and $u_{2}$ solve $\mathcal{L}u=0$ in $\Omega_{\epsilon}$, the comparison principle implies
$\sup_{\Omega_{\epsilon}}|u_{1}-u_{2}|\leq(C_{+}-C_{-})\epsilon^{-\gamma}.$
However, this bound must also hold as $\epsilon\rightarrow 0$. By the precise asymptotic matching at the boundary, the ratio $u_{1}/u_{2}\rightarrow 1$ as $d(x)\rightarrow 0$, which forces $u_{1}=u_{2}$ throughout $\Omega$.
More precisely, consider $w=u_{1}-u_{2}$. Then $w$ satisfies a linear elliptic equation of the form
$\Delta w+c(x)\cdot\nabla w+e(x)w=0$
where $c$ and $e$ are bounded in $\Omega_{\epsilon}$. Since $w=o(d^{-\gamma})$ as $d\rightarrow 0$ (by the sharp asymptotic), the maximum principle applied on $\Omega_{\epsilon}$ with vanishing boundary data in the limit yields $w\equiv 0$.
@@
Set
$u(x)=\sup_{w\in\mathcal{S}}w(x).$
The solution $u$ belongs to $C^{2,\alpha}(\Omega)$ for some $\alpha\in(0,1)$. If additionally $f\in C^{k,\alpha}(\overline{\Omega})$ and $a,b\in C^{k,\alpha}(\Omega)$ for $k\geq 1$, then $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13].
Step 5: Strict convexity.
Let $u$ be the solution obtained. If $\Omega$ is strictly convex, then $u$ is strictly convex in $\Omega$: $D^{2}u(x)>0$ for all $x\in\Omega$.
To prove this, we employ the microscopic convexity principle following Kennington [15] and Korevaar [16].
Step 5.1: Setup.
Define the minimum eigenvalue function
$\lambda(x)=\min_{|\xi|=1}D^{2}u(x)\xi\cdot\xi.$
@@
Step 5.2: Differential inequality.
Suppose $\lambda$ achieves a non-positive minimum at some $x_{0}\in\Omega$. At $x_{0}$, let $\xi_{0}$ be the corresponding eigenvector. Differentiating the equation $\mathcal{L}u=0$ twice in the direction $\xi_{0}$, we obtain (after lengthy but standard calculations):
$\Delta\lambda+c(x)\cdot\nabla\lambda+e(x)\lambda\leq 0\quad\text{at }x_{0},$ (3.3)
where $c$ and $e$ are bounded functions depending on $u$, $\nabla u$, and $D^{2}u$. The key observation is that the strict convexity of $h$ contributes a favorable (non-negative) term to $e(x)$.
Step 5.3: Boundary behavior.
Near $\partial\Omega$, the strict local behavior of $u$ mirrors the explicit analytical structure of the rigorous barrier $W=Cv^{-\gamma}$. Evaluating the exact Hessian of $W$ yields:
Since $v$ is strictly concave ($D^{2}v<0$), the term $C\gamma v^{-\gamma-1}D^{2}v$ is positive definite. The rank-one term $\nabla v\otimes\nabla v$ is positive semi-definite. Therefore, $D^{2}W>0$ near $\partial\Omega$.
By the comparison principle for the Hessian, the boundary behavior forces $D^{2}u\geq D^{2}W>0$ near $\partial\Omega$.
Step 5.4: Strong maximum principle.
Suppose $\lambda(x_{0})=\min_{\overline{\Omega}}\lambda\leq 0$ for some $x_{0}\in\Omega$. By Step 3, $x_{0}$ cannot be near $\partial\Omega$. By the differential inequality (3.3) and the strong maximum principle, either $\lambda\equiv\lambda(x_{0})$ (constant) or the minimum cannot be achieved in the interior.
If $\lambda$ were constant and non-positive, this would contradict the positive definiteness of $D^{2}u$ near $\partial\Omega$ established in Step 3. Therefore, $\lambda>0$ throughout $\Omega$. ∎

### 22 Mar 2026 v2

@@ 3.1. The Defining Function
The strict convexity of $\Omega$ (Assumption 2.1) implies the existence of a smooth strictly convex defining function. Following the approach of Gilbarg and Trudinger [13] and Lazer and McKenna [19], we establish:
@@
Under Assumption 2.1, there exists $\varphi\in C^{2}(\overline{\Omega})$ satisfying:
@@
    $D^{2}\varphi(x)\geq\kappa_{0}\,I$ for all $x\in\overline{\Omega}$, where $\kappa_{0}>0$ is the minimum principal curvature of $\partial\Omega$ from (2.1).
@@
The proof proceeds in two steps.
Step 1: Construction near the boundary. For $x$ sufficiently close to $\partial\Omega$, define $\varphi(x)=-d(x)$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. The signed distance function is $C^{2}$ in a tubular neighborhood $\mathcal{T}_{\delta}=\{x\in\Omega:d(x)<\delta\}$ when the boundary is $C^{2,\alpha}$ (see [13, Chapter 14]). The width of this neighborhood is at least $\delta=1/\kappa_{0}$, where $\kappa_{0}$ is the maximum of the principal curvatures.
The Hessian of d-d at a point $x\in\mathcal{T}_{\delta}$ is related to the second fundamental form of the level set $\{d=\text{const}\}$. By the strict convexity of $\Omega$ (Assumption 2.1), the principal curvatures of these level sets are positive, which gives $D^{2}(-d)(x)\geq\kappa_{0}\,I$ in $\mathcal{T}_{\delta}$.
Step 2: Extension to the interior. The function $\varphi=-d$ can be extended to all of $\overline{\Omega}$ while preserving strict convexity. This is achieved by considering the convex envelope of d-d in the interior region $\Omega\setminus\mathcal{T}_{\delta/2}$ and smoothing via convolution. The detailed construction is given in [13, Chapter 14]. ∎
Let $v(x)=-\varphi(x)$. Then $v>0$ in $\Omega$, $v=0$ on $\partial\Omega$, and $v$ is strictly concave ($D^{2}v<0$). By construction, $v(x)=d(x)$ identically in a tubular neighborhood of $\partial\Omega$, where $|\nabla v(x)|=1$ holds. These properties of $v$ are used extensively in the barrier construction that follows.
@@ 3.2. Explicit Barrier Construction
We now construct explicit sub- and supersolutions that determine the precise boundary blow-up rate. This construction is the technical core of the paper.
@@
    Critical Logarithmic Case ($q=\beta+2$): The singular growth is logarithmic, $u(x)\sim C\,\ln(1/v(x))$. There exist $C_{+}>C_{-}^{*}>0$ and $0<C_{-}<C_{-}^{*}$ providing local barriers near the boundary.
@@
We provide detailed computations for each case, using all hypotheses explicitly.
@@
We compute the gradient, its norm, and the Laplacian of $W$, using only the chain rule and the properties of $v$ from Lemma 3.1:
@@
For the Hessian of $W$, we compute:
Step 2: Substitution into the operator $\mathcal{L}$.
Substituting (3.3)–(3.5) into $\mathcal{L}W$ defined in (3.1):
@@
$\quad+b(x)\,h\bigl(C\,\gamma\,v^{-\gamma-1}\,|\nabla v|\bigr)+a(x)\,C\,v^{-\gamma}-f(x).$ (3.7)
@@
We now use the weight hypotheses (W1)–(W2) from Assumption 2.6 and the Hamiltonian bounds (H3) from Assumption 2.4. Near $\partial\Omega$, the properties of $v$ from Lemma 3.1 give $v(x)=d(x)$ and $|\nabla v(x)|=1$.
For any $\varepsilon>0$, by (2.4), there exists $\delta>0$ such that for all $x\in\Omega$ with $v(x)<\delta$:
$(b_{1}-\varepsilon)\,v(x)^{\beta}\leq b(x)\leq(b_{2}+\varepsilon)\,v(x)^{\beta}.$ (3.8)
Using the lower bound in (H3), the gradient term satisfies:
@@
$=(b_{1}-\varepsilon)\,l_{1}\,(C\,\gamma)^{q}\,v^{\beta-q(\gamma+1)}.$ (3.9)
The leading singular terms in $\mathcal{L}W$ are:
@@
where $\mathcal{R}(x)$ denotes terms of strictly lower singular order relative to $v^{-\gamma-2}$, specifically:
$\mathcal{R}(x)=C\,\gamma\,v^{-\gamma-1}\,\Delta v+a(x)\,C\,v^{-\gamma}-f(x).$ (3.11)
Step 4: Balance of leading singularities.
We choose $\gamma$ to perform an exact balance of the leading singularities by requiring the exponent of $v$ inside the brackets in (3.10) to vanish:
$\beta-q(\gamma+1)+\gamma+2=0\iff\gamma(1-q)=\beta-q+2\iff\gamma=\frac{\beta-q+2}{q-1}.$ (3.12)
Note that $\gamma>0$ if and only if $\beta-q+2>0$, i.e., $q<\beta+2$. This is the Gradient-Dominant regime.
With this choice of $\gamma$, the expression in brackets becomes:
$P(C)=(b_{1}-\varepsilon)\,l_{1}\,\gamma^{q}\,C^{q}-\gamma\,(\gamma+1)\,C.$ (3.13)
@@
  • $P(C)<0$ for small $C>0$ (the linear term $\gamma(\gamma+1)C$ dominates);
  • $P(C)>0$ for large $C$ (the power term $C^{q}$ dominates);
  • $P(C^{*})=0$ at the critical constant $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}\,l_{1}\,\gamma^{q}}\right)^{1/(q-1)}$.
@@
Step 5: Verification of the subordinate terms.
The reaction term, using (W1) from Assumption 2.6:
$a(x)\,C\,v^{-\gamma}\leq a_{2}\,C\,v^{\alpha-\gamma}.$ (3.14)
Since $\alpha>-2$ by (W1), we have $\alpha-\gamma>-2-\gamma=-\gamma-2$, so the reaction term is of order $v^{\alpha-\gamma}$, which is strictly lower than the leading order $v^{-\gamma-2}$. Therefore, after multiplying by $v^{\gamma+2}$, this term vanishes in the limit $v\to 0$.
@@
The source term $f(x)$ is bounded on $\overline{\Omega}$ by Assumption 2.3, hence $f(x)=O(1)=o(v^{-\gamma-2})$.
Step 6: Case 2 — High-order gradient.
When $q>\beta+2$, equation (3.12) gives $\gamma<0$ (the numerator $\beta-q+2$ is negative). Thus, for any chosen $\gamma>0$, the exponent $\beta-q(\gamma+1)+\gamma+2<0$, meaning $v^{\beta-q(\gamma+1)}\cdot v^{\gamma+2}$ has a negative exponent. The gradient term is therefore more singular than the Laplacian term, and $\mathcal{L}W>0$ near $\partial\Omega$ for any $C>0$.
Step 7: Case 3 — Critical logarithmic.
When $q=\beta+2$, let $W=C\,\ln(1/v)=-C\,\ln v$. Then:
$\nabla W$ $=-C\,v^{-1}\,\nabla v,$ (3.15)
$|\nabla W|$ $=C\,v^{-1}\,|\nabla v|,$ (3.16)
$\Delta W$ $=C\,v^{-2}\,|\nabla v|^{2}-C\,v^{-1}\,\Delta v.$ (3.17)
Substituting into $\mathcal{L}W$ and using $|\nabla v|=1$, $b(x)\geq(b_{1}-\varepsilon)\,v^{\beta}$, $h(s)\geq l_{1}\,s^{q}$:
$\mathcal{L}W$ $\geq-C\,v^{-2}+C\,v^{-1}\,\Delta v+(b_{1}-\varepsilon)\,v^{\beta}\,l_{1}\,(C\,v^{-1})^{q}+a_{1}\,v^{\alpha}\,C\,\ln(1/v)-f$
$=v^{-2}\left[(b_{1}-\varepsilon)\,l_{1}\,C^{q}\,v^{\beta-q+2}-C\right]+\mathcal{R}_{L}(x),$ (3.18)
where $\mathcal{R}_{L}(x)$ denotes terms of strictly lower singular order. Since $\beta-q+2=0$ in the critical case, we obtain:
$\liminf_{x\to\partial\Omega}\left(v^{2}\,\mathcal{L}W\right)\geq(b_{1}-\varepsilon)\,l_{1}\,C^{q}-C.$ (3.19)
The same argument as in Step 4 shows that barriers exist with the appropriate choice of $C_{\pm}$. ∎
@@
The strict convexity of $\Omega$ is essential in two ways: (i) it guarantees the existence of a smooth, strictly concave defining function $v$ (Lemma 3.1), ensuring $D^{2}v<0$; (ii) it ensures that $|\nabla v|$ is bounded away from zero near $\partial\Omega$, which is necessary for the gradient term to provide the required singular behavior. Without strict convexity, the gradient could vanish at flat boundary points, destroying the singularity balance.
@@ 4. Existence, Uniqueness, and Convexity
$u$ $\leq w\quad\text{on }\partial\Omega^{\prime}.$ (4.1)
@@
for some $x_{0}\in\Omega^{\prime}$. The maximum cannot be achieved on $\partial\Omega^{\prime}$ by the boundary hypothesis (4.1). At the interior maximum point $x_{0}$:
$\nabla u(x_{0})$ $=\nabla w(x_{0}),$ (4.2)
$\Delta u(x_{0})$ $\leq\Delta w(x_{0}).$ (4.3)
Using (4.2), the Hamiltonian terms coincide: $h(|\nabla u(x_{0})|)=h(|\nabla w(x_{0})|)$. Therefore:
@@
$=-\Delta u(x_{0})+\Delta w(x_{0})+b(x_{0})\bigl[h(|\nabla u(x_{0})|)-h(|\nabla w(x_{0})|)\bigr]$
$\quad+a(x_{0})\bigl[u(x_{0})-w(x_{0})\bigr]$
$\geq a(x_{0})\bigl[u(x_{0})-w(x_{0})\bigr]>0,$
where we used (4.3) for the first term, (4.2) for the vanishing of the Hamiltonian difference, and $a(x_{0})>0$ from Assumption 2.6. This is a contradiction. ∎
@@
Let $\overline{u}(x)=C_{+}\,v(x)^{-\gamma}$ and $\underline{u}(x)=C_{-}\,v(x)^{-\gamma}$ be the super- and subsolutions from Lemma 3.2. By construction, these satisfy
@@
for some $\delta_{0}>0$ determined by the asymptotic estimates in Lemma 3.2.
For the interior region $\Omega_{\delta_{0}}=\{x\in\Omega:d(x)\geq\delta_{0}\}$, we solve the Dirichlet problem
@@
Similarly construct a global subsolution $\underline{U}$ with $\underline{U}\leq\overline{U}$ throughout $\Omega$.
@@
$\mathcal{S}=\bigl\{w\in C(\Omega):w\text{ is a subsolution},\;\underline{U}\leq w\leq\overline{U}\bigr\}.$
@@
Taking the supremum over all such modifications, standard arguments [10] show that $u$ is a viscosity solution. By interior regularity for uniformly elliptic equations with Hölder continuous coefficients, $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$, and $u$ is a classical solution. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13].
@@
Suppose $u_{1}$ and $u_{2}$ are two solutions satisfying (2.5). Consider $w=u_{1}-u_{2}$. Then $w$ satisfies a linearized equation of the form
$\Delta w+c(x)\cdot\nabla w+e(x)\,w=0,$ (4.6)
where the coefficients $c(x)$ and $e(x)$ are obtained from the mean value theorem applied to $h(|\nabla u|)$ and a(x)ua(x)\,u:
$c(x)$ $=b(x)\,\frac{h(|\nabla u_{1}|)-h(|\nabla u_{2}|)}{|\nabla u_{1}|-|\nabla u_{2}|}\,\frac{\nabla(|\nabla u_{1}|-|\nabla u_{2}|)}{|\nabla u_{1}|-|\nabla u_{2}|},$
$e(x)$ $=a(x).$
Since $w=u_{1}-u_{2}=o(d^{-\gamma})$ as $d\to 0$ (because both $u_{1}$ and $u_{2}$ satisfy the same sharp asymptotics (2.5)), the function $w$ grows slower than the barriers near the boundary.
For any $\epsilon>0$, define $\Omega_{\epsilon}=\{x\in\Omega:d(x)>\epsilon\}$. On $\partial\Omega_{\epsilon}$, both $u_{1}$ and $u_{2}$ are finite. The maximum principle applied to (4.6) on $\Omega_{\epsilon}$ gives:
$\sup_{\Omega_{\epsilon}}|w|\leq\sup_{\partial\Omega_{\epsilon}}|w|.$
Since $|w|=|u_{1}-u_{2}|=o(d^{-\gamma})$ on $\partial\Omega_{\epsilon}$ (where $d=\epsilon$), and the precise asymptotic matching forces $u_{1}/u_{2}\to 1$ as $d\to 0$, we conclude that $w\equiv 0$ by taking $\epsilon\to 0$.
@@
The solution $u$ defined by (4.5) belongs to $C^{2,\alpha}(\Omega)$ for some $\alpha\in(0,1)$. If additionally $f\in C^{k,\alpha}(\overline{\Omega})$ and $a,b\in C^{k,\alpha}(\Omega)$ for $k\geq 1$, then $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13]. ∎
Proof of Theorem 2.11.
The proof of strict convexity proceeds in four steps, using the microscopic convexity principle following Kennington [15] and Korevaar [16].
Step 1: Setup.
Let $u$ be the solution from Theorem 2.9. Define the minimum eigenvalue function
$\lambda(x)=\min_{|\xi|=1}D^{2}u(x)\,\xi\cdot\xi.$ (4.7)
@@
Step 2: Differential inequality.
Suppose $\lambda$ achieves a non-positive minimum at some $x_{0}\in\Omega$. At $x_{0}$, let $\xi_{0}$ be the corresponding eigenvector. Differentiating the equation $\mathcal{L}u=0$ twice in the direction $\xi_{0}$, and using the fact that $h$ is strictly convex (Assumption 2.4, condition (H2)), we obtain:
$\Delta\lambda+c(x)\cdot\nabla\lambda+e(x)\,\lambda\leq\mathcal{Q}(x,u,\nabla u)\quad\text{at }x_{0},$ (4.8)
where $c$ and $e$ are bounded functions, and the quadratic form $\mathcal{Q}$ is given by:
$\mathcal{Q}=-\bigl[b^{\prime\prime}(x)\,h(|\nabla u|)+a^{\prime\prime}(x)\,u-f^{\prime\prime}(x)\bigr],$ (4.9)
where $b^{\prime\prime}$, $a^{\prime\prime}$, $f^{\prime\prime}$ denote second directional derivatives in the direction $\xi_{0}$.
Under the hypotheses that $a(x)$ and $b(x)$ are convex (Assumption 2.6) and $f(x)$ is concave (Assumption 2.3), we have $a^{\prime\prime}\geq 0$, $b^{\prime\prime}\geq 0$, and $f^{\prime\prime}\leq 0$. Since $h\geq 0$, $u>0$, and $f\geq 0$, we conclude:
$\mathcal{Q}(x,u,\nabla u)\leq 0.$ (4.10)
Step 3: Boundary behavior.
Near $\partial\Omega$, the solution $u$ is well-approximated by the barrier $W=C\,v^{-\gamma}$. The Hessian of $W$, computed in (3.6), is:
$D^{2}W=C\,\gamma\,(\gamma+1)\,v^{-\gamma-2}\,\nabla v\otimes\nabla v-C\,\gamma\,v^{-\gamma-1}\,D^{2}v.$
Since $v$ is strictly concave ($D^{2}v<0$ by Lemma 3.1), the term $C\,\gamma\,v^{-\gamma-1}\,D^{2}v$ is positive definite. The rank-one term $\nabla v\otimes\nabla v$ is positive semi-definite. Therefore, $D^{2}W>0$ near $\partial\Omega$, which by comparison forces $D^{2}u>0$ near the boundary.
Step 4: Strong maximum principle.
Suppose $\lambda(x_{0})=\min_{\overline{\Omega}}\lambda\leq 0$ for some $x_{0}\in\Omega$. By Step 3, $x_{0}$ cannot be near $\partial\Omega$. By the differential inequality (4.8) with the favorable sign (4.10), and the strong maximum principle for linear elliptic equations, either $\lambda\equiv\lambda(x_{0})$ (constant) or the minimum cannot be achieved in the interior.
If $\lambda$ were constant and non-positive, this would contradict the positive definiteness of $D^{2}u$ near $\partial\Omega$ established in Step 3. Therefore, $\lambda>0$ throughout $\Omega$, which means $D^{2}u(x)>0$ for all $x\in\Omega$. ∎